# How to Correctly Report LLM-as-a-Judge Evaluations

### v2

@@ 7 Robust Estimation under Distribution Shift
Table 1: Comparison of point estimators for estimating $\mathbb{E}_{{\mathbb{P}}}[{\mathbf{Z}}]$, together with their estimation formulas and assumptions for unbiasedness. Here, $\mathbb{P}$} denotes the test distribution and $\mathbb{Q}$} the calibration distribution. Estimators (ii)–(iv) require $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$ for their assumptions to hold, whereas the misclassification-adjusted estimator in (v), used in our method, remains valid when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$.
@@
(v) Misclassification-adjusted estimator (Ours; extending rogan1978estimating) $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{\!-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$
@@
LLM-as-a-judge evaluation can be formalized through a data-generating process consisting of two probabilistic components: the marginal distribution of true labels $\mathop{\rm Pr}\nolimits(Z)$, which is determined by the dataset, and the conditional behavior of the LLM judge, $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z,\xi)$. Here, $\xi$ denotes auxiliary factors that may influence judgments, such as response length or stylistic attributes (zhou2023instruction; dubois2024lengthcontrolled).
Throughout the main analysis, we only assume that the judge’s behavior depends only on the true label and is invariant across datasets:
@@
This assumption reflects the common practice of using the same LLM judge and prompting strategy for both test and calibration datasets. We discuss a relaxation of this assumption that allows $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z,\xi)$ to depend on $\xi$ in Section 9.
In contrast, we allow the label distributions to differ:
@@
Such distribution shift can arise because calibration datasets are curated to facilitate accurate error estimation (e.g., through class balancing or simplified examples), whereas test datasets are designed to reflect realistic evaluation scenarios (jung2024trust).
@@
We compare several point estimators for estimating $\mathbb{E}_{{\mathbb{P}}}[Z]$ under the test distribution, as studied in prior work (kloos2021comparing; meertens2022improving). To facilitate comparison, define
@@
Table 1 summarizes these estimators, along with their formulas and the invariance assumptions required for unbiasedness. As discussed earlier, the naive estimator in (i), $\hat{p}=[\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}]]_{1}$, is biased when the LLM judge is imperfect.
Figure 5: Effects of distribution shift $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$ on estimator bias. We fix the true test-set accuracy at $\theta_{{\mathbb{P}}}=0.5$ and vary the calibration-set accuracy $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, corresponding to a distribution-shift ratio from $0.5$ to $1.5$. The misclassification-adjusted estimator in (v), used in our method, remains unbiased.
(a) Estimates
(b) Coverage probability
(c) Confidence interval length
Figure 6: Monte Carlo simulation for estimating $\theta$ under an imperfect LLM judge with $(q_{0},q_{1})=(0.7,0.9)$. We evaluate estimators across 21 values of $\theta\in[0,1]$, each visualized as a single point. Figure 6(a) reports the results from a single run, while Figure 6(b) and Figure 6(c) summarize averages computed over 10,000 replications. All experiments use a test dataset of size $n=1000$ and a calibration dataset of size $m=200$, and we use an equal allocation $m_{0}=m_{1}$ for Figure 6(a) and Figure 6(c). (a) The naive estimator $\hat{p}$ in Eq. (3) exhibits bias, while the unbiased estimator $\hat{\theta}$ in Eq. (8) closely recovers the true accuracy $\theta$ across all values. Shaded regions represent the 95% confidence intervals (CI). (b) Across all $\theta$, the coverage probability of the confidence interval remains consistently close to the nominal 95% level. (c) Given a fixed calibration budget of $m=200$, we compare two allocation strategies: an equal split ($m_{0}=m_{1}$) and the allocation proportional to $m_{0}\propto(1/\hat{p}-1)\sqrt{\kappa}\cdot m_{1}$ by using Algorithm 1. The proposed allocation gives shorter confidence intervals.
@@
Most existing calibration-based estimators rely on invariance assumptions that fail under the distribution shift described in Eq. (21), even if the judge’s conditional behavior remains stable. In particular, the calibration-only estimator in (ii), the difference estimator in (iii), and the conditional calibration estimator in (iv) all require invariance of the marginal label distribution $\mathop{\rm Pr}\nolimits(Z)$. This assumption is violated under the distribution shift.
In contrast, the misclassification-adjusted estimator in (v), used in our method, requires only stability of the confusion matrix $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z)$ across datasets. Because it does not depend on the label distribution $\mathop{\rm Pr}\nolimits(Z)$, it remains unbiased even when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$. In Appendix B.3, we show that this estimator, used in our method, is equivalent to our point estimator defined in Eq. (8).
To illustrate this robustness, consider a setting in which the test and calibration datasets share the same conditional judge behavior with $(q_{0},q_{1})=(0.7,0.9)$, but differ in their true accuracies: $\theta_{{\mathbb{P}}}=0.5$ for the test dataset and $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$ for the calibration dataset. The resulting biases are shown in Figure 5. Under distribution shift, the calibration-only estimator (ii), difference estimator (iii) and the difference estimator (iv) exhibit bias, whereas the misclassification-adjusted estimator (v) remains unbiased, as it does not depend on the marginal distribution of $Z$. In other words, it remains valid even when $\theta_{{\mathbb{P}}}\neq\theta_{{\mathbb{Q}}}$. Details of the simulation are provided in Appendix D.
8 Empirical Validation on Our Method
To validate the theoretical results established in Section 5, we empirically evaluate the proposed estimator, confidence interval, and calibration allocation strategy through Monte Carlo simulation and real-world benchmarks.
8.1 Monte Carlo Simulation
Experimental Setup.
We evaluate the proposed method under the following parameter configuration. The LLM judge is characterized by parameters $(q_{0},q_{1})=(0.7,0.9)$, and the true accuracy varies over $\theta\in\{0,0.05,0.10,\ldots,1\}$, resulting in 21 distinct settings. For each configuration of $(q_{0},q_{1},\theta)$, we generate a test dataset of size $n=1000$ and a calibration dataset of total size $m=200$, with equal allocation $m_{0}=m_{1}$ unless stated otherwise.
We compute the naive estimator $\hat{p}$ in (3) together with its confidence interval, and compute the bias-adjusted estimator $\hat{\theta}$ in (8) and the confidence interval in (9). Each configuration is replicated 10,000 times to evaluate estimates, interval coverage, and interval length. Additional simulations under alternative settings are provided in Appendix E.
Bias Reduction in Point Estimation.
Figure 6(a) compares the naive estimator $\hat{p}$ and the bias-adjusted estimator $\hat{\theta}$ based on a single simulation run. As shown in Section 5, $\hat{p}$ exhibits bias, particularly overestimating the true accuracy when the underlying $\theta$ is small. In contrast, $\hat{\theta}$ closely aligns with the true accuracy across all values of $\theta$, demonstrating the bias correction achieved by (8).
Coverage of the Confidence Interval.
Figure 6(b) reports the empirical coverage probability of the confidence interval in (9). Across all values of $\theta$, the coverage remains consistently close to the nominal 95% level, whereas the confidence interval constructed from the naive estimator $\hat{p}$ achieves nearly zero coverage except at a few values of $\theta$. These results confirm that the proposed confidence interval provides reliable uncertainty quantification.
Efficiency of Optimal Calibration Allocation.
To examine the benefits of optimal calibration allocation, we compare two strategies under a fixed calibration budget of $m=200$: (i) an equal allocation ($m_{0}=m_{1}=100$), and (ii) the allocation produced by Algorithm 1, which approximates the optimal ratio derived in Proposition 5.2. Figure 6(c) shows that the adaptive allocation consistently gives shorter confidence intervals than the equal-split baseline.
8.2 Chatbot Arena Benchmark
We evaluate our method on the Chatbot Arena benchmark (chiang2024chatbot), a crowdsourced platform widely used in LLM-as-a-judge studies (zheng2023judging). In Chatbot Arena, users vote between responses generated by two anonymous models to the same prompt. These pairwise comparisons are aggregated to estimate each model’s win rate, defined as the probability that the target model’s response is preferred over its opponent’s.
Experimental Setup.
We conduct experiments on six models used in zheng2023judging: Alpaca-13B (alpaca), Claude-v1, FastChat-T5-3B, GPT-4, LLaMA-13B (touvron2023llama), and Vicuna-13B (vicuna2023). For each target model, the dataset consists of response pairs composed of one response generated by the target model and one generated by a randomly selected opponent model.
To quantify bias, we consider question–response pairs with human preference evaluations, which serve as the ground truth for bias computation. Each response pair is additionally evaluated by GPT-4.1-mini, which acts as the LLM judge whose bias we aim to analyze and correct.
The dataset is randomly split into a test set (90%) and a calibration set (10%). The calibration set is used to estimate parameters for bias adjustment, while bias is evaluated on the test set. We repeat this process 100 times with different random splits and report averaged results.
We compare two estimators: the baseline estimator $\hat{p}$, which directly uses the LLM judge’s preference probabilities, and our $\hat{\theta}$ which applies the proposed bias-adjustment method.
Figure 7: Average bias of winning rates for six models on Chatbot Arena, averaged over 100 random test (90%) / calibration (10%) splits. We compare the raw LLM judge estimator $\hat{p}$ and the proposed bias-adjusted estimator $\hat{\theta}$, using GPT-4.1-mini as the LLM judge. The proposed method reduces bias across all models.
Table 2: 95% confidence interval coverage and average CI length of winning rate estimates on Chatbot Arena, averaged over 100 random replications.
Model Alpaca Claude FastChat-T5 GPT-4 LLaMA Vicuna
Coverage 96% 95% 97% 97% 99% 99%
CI length 0.193 0.262 0.256 0.210 0.267 0.178
Results.
Figure 7 shows the average bias of winning rate for each model on the Chatbot Arena benchmark. Using raw judge scores $\hat{p}$, all models exhibit non-trivial bias with varying magnitude and direction. In contrast, the proposed bias-adjustment method $\hat{\theta}$ reduces bias across all six models. The improvement is consistent, indicating effective mitigation of judge-induced bias.
Table 2 reports the empirical coverage rates of the proposed 95% confidence intervals. Across all models, the coverage rates are close to or slightly above the nominal level, indicating that the proposed intervals preserve the desired significance level.
Overall, our method provides a simple and practical solution for bias correction in LLM-based evaluation. It operates directly on existing judge scores and requires only a small calibration dataset, enabling a principled statistical treatment of LLM-as-a-judge evaluation pipelines.
@@ 9 Conclusion
In LLM-as-a-judge evaluation, noisy judgments induce bias in naive point estimates. We introduce a bias-adjusted estimator that corrects for imperfect judgments and constructs confidence intervals that account for uncertainty arising in both the evaluation and calibration datasets. To reduce uncertainty, we show that focusing on the calibration design, such as the allocation between response types, is preferable to simply increasing the total number of judgments.
Beyond bias correction, we demonstrate that when the LLM judge is sufficiently accurate, our framework can achieve lower variance than direct human evaluation, and that the resulting estimator remains robust to distribution shift between the test and calibration datasets under a mild and realistic assumption on the LLM judge. We hope this work contributes to more reliable and transparent reporting practices in LLM-based evaluation.
@@
Several directions remain for future work. (1) Our method can be extended to account for auxiliary factors that influence LLM-as-a-judge evaluations beyond the true label $\mathbf{Z}$}. While our current analysis assumes that the LLM evaluation $\hat{{\mathbf{Z}}}$ depends only on the true label $\mathbf{Z}$}, in practice it may also be affected by additional nuisance factors $\xi$, such as response length or other stylistic attributes. This more realistic setting can be accommodated by allowing both the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ and the LLM evaluation $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ to depend on such factors. For example, one may estimate separate confusion matrices across different strata of $\xi$ (e.g., short vs. long responses) and perform bias adjustment within each group, followed by aggregation across strata. Formally, this corresponds to correcting $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ using $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ at the stratum level and averaging over $\xi$. (2) The proposed method can be extended to a multinomial setting by generalizing $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Sec. 7 to a multivariate formulation, where $\mathbf{Z}$} and $\hat{{\mathbf{Z}}}$ are represented as probability distributions over multiple response categories. Such an extension would increase the number of parameters, and additional structural assumptions, such as constraints on the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$, may be required to ensure stable estimation. (3) A conformal prediction framework could be incorporated to provide sample-specific uncertainty quantification by constructing statistically valid prediction regions around each bias-adjusted point estimate.

### v3

@@ 7.2 Chatbot Arena Benchmark
Table 2: Existing bias-correction estimators for $\mathbb{E}_{{\mathbb{P}}}[{\mathbf{Z}}]$. Here, $\mathbb{P}$} denotes the test distribution and $\mathbb{Q}$} the calibration distribution. Estimators (ii)–(iv) rely on the assumption that $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}({\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}({\mathbf{Z}})$, whereas our estimator in (v) remains valid when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}({\mathbf{Z}})\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}({\mathbf{Z}})$.
@@
(v) Misclassification-adjusted estimator (Ours; $\hat{\theta}$ in (7)) $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{\!-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$
Results.
Figure 6 shows the average bias in win-rate estimation for each model on the Chatbot Arena benchmark. The naive estimator $\hat{p}$ exhibits non-negligible bias across all six models, whereas our estimator $\hat{\theta}$ reduces bias toward zero. Table 1 reports the coverage of the proposed 95% confidence intervals, which remains close to or slightly above the nominal level across models, indicating valid uncertainty quantification. Overall, our method provides a simple and practical bias-correction procedure for LLM-as-a-judge evaluation that requires only a small calibration dataset.
8 Robust Estimation under Distribution Shift
As mentioned in Section 1, a variety of bias-correction methods have been proposed for imperfect evaluators, typically under the assumption $\mathbb{P}$$\mathbb{Q}$} between the test distribution $\mathbb{P}$} and the calibration distribution $\mathbb{Q}$}. When this assumption is violated, bias correction can fail.
In practice, however, not every marginal or conditional probability under $\mathbb{P}$} need coincide with its counterpart under $\mathbb{Q}$}. Instead, only certain conditional distributions of the data-generating process may remain invariant across $\mathbb{P}$} and $\mathbb{Q}$}, while others may change. We therefore clarify which probabilistic assumptions are plausibly shared between $\mathbb{P}$} and $\mathbb{Q}$} in practice. Under minimal assumptions that are realistic in LLM-as-a-judge evaluation, we then characterize which methods remain unbiased under distribution shift.
@@ 8 Robust Estimation under Distribution Shift
LLM-as-a-judge evaluation can be formalized through a data-generating process consisting of two probabilistic components: the marginal distribution of true labels $\mathop{\rm Pr}\nolimits(Z)$, which is determined by the dataset, and the conditional behavior of the LLM judge, $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z,\xi)$. Here, $\xi$ denotes auxiliary factors that may influence judgments, such as response length or stylistic attributes (Zhou et al., 2023; Dubois et al., 2024).
In this section, we only assume that the judge’s behavior depends on the true label and that it remains invariant across the test and calibration datasets:
@@
This assumption is motivated by common practice in LLM-as-a-judge evaluation, where the same LLM judge and prompting strategy are used for both the test and calibration datasets. In Appendix 9, we discuss how this assumption can be relaxed to allow the judge behavior to depend on additional covariates $\xi$, namely $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(\hat{Z}\mid Z,\xi)=\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{Z}\mid Z,\xi)$.
In contrast, we consider the true label distributions to differ:
@@
Such distribution shift can arise because datasets are collected in different ways: calibration datasets are constructed to analyze how the judge makes errors, whereas test datasets reflect realistic evaluation scenarios (Jung et al., 2024). Since calibration datasets require costly human annotation, they are collected in advance and remain fixed over time. By contrast, test data are often collected online or arrive in multiple batches, each potentially having a different underlying true accuracy. Consequently, the true label distribution of the test data can differ from that of the calibration data.
@@
Following prior work (Kloos et al., 2021; Meertens et al., 2022), we compare five point estimators for estimating the true accuracy $\theta$ in (2) under the test distribution, that is, $\mathbb{E}_{{\mathbb{P}}}[Z]$. To facilitate comparison, we define
@@
[…]
@@
Most existing estimators rely on invariance assumptions that fail under the distribution shift in true accuracy described in (19), even when the judge’s conditional behavior remains unchanged. In particular, the estimators in (ii)–(iv) require invariance of the marginal distribution $\mathop{\rm Pr}\nolimits(Z)$, which is violated when $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$. In contrast, the misclassification-adjusted estimator in (v) depends only on the conditional distribution $\mathop{\rm Pr}\nolimits(\hat{Z}\mid Z)$ and does not rely on the marginal distribution $\mathop{\rm Pr}\nolimits(Z)$.
Following the Monte Carlo simulation in Section 7, we examine estimator behavior under distribution shift. We consider a setting where the test and calibration datasets share the same conditional judge behavior, while the calibration accuracy varies as $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, with the test accuracy held fixed at $\theta_{{\mathbb{P}}}=0.5$.
Figure 7 reports the resulting biases of each estimator. As established in previous sections, the naive estimator in (i) is biased in this setting. Estimators (ii)–(iv) also exhibit bias when $\theta_{{\mathbb{P}}}\neq\theta_{{\mathbb{Q}}}$, whereas the misclassification-adjusted estimator in (v), used in our method, remains unbiased. Additional simulation details are provided in Appendix E.
Figure 7: Effects of distribution shift $\mathop{\rm Pr}\nolimits_{{\mathbb{P}}}(Z)\neq\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(Z)$ on estimator bias. We fix the true accuracy of the test distribution at $\theta_{{\mathbb{P}}}=0.5$ and vary the true accuracy of the calibration distribution as $\theta_{{\mathbb{Q}}}\in[0.25,0.75]$, corresponding to a distribution-shift ratio from $0.5$ to $1.5$. The misclassification-adjusted estimator in (v), used in our method, remains unbiased under this shift.
@@ 9 Conclusion
In LLM-as-a-judge evaluation, noisy judgments can induce bias in naive estimates. We thus introduce a bias-corrected estimator and provide statistically sound confidence intervals. To further reduce interval length, we show that calibration design using the proposed algorithm, which allocates samples across response types, can be an effective strategy.
Beyond bias correction, we show that when the LLM judge is sufficiently accurate, our framework can achieve lower variance than human-only evaluation. Moreover, under a mild and realistic assumption on the LLM judge, our estimator remains unbiased and robust to distribution shift between the test and calibration datasets. Several directions for future work are discussed in Appendix 9. We hope this work contributes to more reliable, interpretable, and transparent practices for LLM-based evaluation.
@@
Several directions remain for future work. (1) Our method can be extended to account for auxiliary factors that influence LLM-as-a-judge evaluations beyond the true label $\mathbf{Z}$}. While our current analysis assumes that the LLM evaluation $\hat{{\mathbf{Z}}}$ depends only on the true label $\mathbf{Z}$}, in practice it may also be affected by additional nuisance factors $\xi$, such as response length or other stylistic attributes. This more realistic setting can be accommodated by allowing both the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ and the LLM evaluation $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ to depend on such factors. For example, one may estimate separate confusion matrices across different strata of $\xi$ (e.g., short vs. long responses) and perform bias adjustment within each group, followed by aggregation across strata. Formally, this corresponds to correcting $\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}\mid\xi]$ using $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}},\xi)$ at the stratum level and averaging over $\xi$. (2) The proposed method can be extended to a multinomial setting by generalizing $\big(\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})\big)^{-1}\displaystyle\mathbb{E}_{{\mathbb{P}}}[\hat{{\mathbf{Z}}}$] in Section 8 to a multivariate formulation, where $\mathbf{Z}$} and $\hat{{\mathbf{Z}}}$ are represented as probability distributions over multiple response categories. Such an extension would increase the number of parameters, and additional structural assumptions, such as constraints on the confusion matrix $\mathop{\rm Pr}\nolimits_{{\mathbb{Q}}}(\hat{{\mathbf{Z}}}\mid{\mathbf{Z}})$, may be required to ensure stable estimation. (3) A conformal prediction framework could be incorporated to provide sample-specific uncertainty quantification by constructing statistically valid prediction regions around each point estimate.
Acknowledgements
This work was supported by the National Science Foundation (NSF) Award DMS-2023239, the NSF CAREER Award CCF-2339978, an Amazon Research Award, and a grant from FuriosaAI. In addition, it was supported by the National Research Foundation of Korea (NRF) grant funded by the Korean Ministry of Science and ICT (MSIT) (RS-2024-00345351, RS-2024-00408003), by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by MSIT (RS-2023-00259934, RS-2025-02283048), and by the Brain Korea 21 program funded by the Korean Ministry of Education and the NRF (BK21 FOUR, No.5199990913980).